OpenStudy (anonymous):

sec-tan=cos/1+sin I have to get the left to become the right without changing the right hand side using trig fomulas

OpenStudy (anonymous):

So how can you rewrite the left-hand side to start out?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No, how would you do it? Just making sure you know how to start, lol.

OpenStudy (anonymous):

I changed the sec and tan to 1/cos and sin/cos

OpenStudy (anonymous):

Right. And both of those fractions have a denominator of cos and can thus be combined into \(\frac{1-sinx}{cosx}\) I'm sure that makes sense. From there, the trick is to multiply top and bottom by the conjugate of the numerator. As in multiply top and bottom by 1+sinx. So what would the numerator become if you wee to do that?

OpenStudy (anonymous):

(1-sin^2)/cos(1+sin)?

OpenStudy (anonymous):

Exactly. And you can use an identity on the \(1-sin^{2}x\)

OpenStudy (anonymous):

cos^2/(1+sin)

OpenStudy (anonymous):

RIght, you would now have \[\frac{ \cos^{2}x }{ cosx(1+sinx) }\] which from there you can seen how one of the cosines would cancel and youd have the result youre looking for :)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem :)